An object is thrown straight down from the top of a building at a speed of 43 m/s. It hits the ground with a speed of 87 m/s. a) How high is the building? b) How long was the object in the air? (note that there is an acceleration due to gravity).


1
Expert's answer
2021-10-10T16:06:43-0400

(a) Write the equation for speed vs. height:


"v_f^2=v_i^2+2gh,\\\space\\
h=\frac{v_f^2-v_i^2}{2g}=292\text{ m}."

(b) The time is


"t=\frac{v_f-v_i}g=4.49\text{ s}."